# Combinatorial Optimization using Comparison Oracles

### v1

@@ A.1 The Certification Problem
Before addressing the optimization problem, we ask whether it is even possible to certify optimality efficiently. Suppose we wish to prove that some point $y\in\mathcal{P}$ is the optimal solution, i.e., $y=\arg\min_{x\in\mathcal{P}}\langle w^{*},x\rangle$, or equivalently, that
@@ A.3 Beyond the Hypercube
Appendix B Proofs from Section˜3
Proof of Lemma˜3.6.
@@ Appendix B Proofs from Section˜3
Appendix C Proofs from Section˜4
@@
Lemma C.2 (tipping point).
@@
By ˜C.1, the $\textnormal{sign}(|\partial(S)|-|\partial(S+u)|)$ is monotonic in $S$, since the quantity $|\partial(u,S)|$ increases as $S$ grows. Therefore, if the sign changes between $S_{0}$ and $S_{t}$, then by monotonicity, it must strictly increase at some index $i$, which proves (12). Moreover, using the equality in (11) again, (12) can be rewritten as
@@
Lemma C.3 (median sets).
@@
Let $L_{j}=\{v_{1},v_{2},\dots\dots,v_{j}\}$ and $R_{j}=\{v_{j},v_{j+1},\dots,v_{n-1}\}$ for $j\in[n-1]$. First, check if $\partial(v_{1})>\partial(v_{1}+u)$ in which case, we can conclude $S_{u}^{-}=\emptyset,S_{u}^{+}=\{v_{2},v_{3},\dots,v_{n-1}\}$ because $\{u,v_{1}\}$ has to be the only edge using ˜C.1. Similarly for $v_{n-1}$. So from now on, assume that $\partial(v_{1})\leq\partial(v_{1}+u)$ and $\partial(v_{n-1})\leq\partial(v_{n-1}+u)$. Equivalently, this can also be written as $\partial(R_{2})\geq\partial(R_{2}+u),\partial(L_{n-2})\geq\partial(L_{n-2}+u)$.
In what follows, we will first describe how to compute $S_{u}^{-}$ and then use a similar argument to compute $S_{u}^{+}$. If $\partial(v_{1})=\partial(v_{1}+u)$, then again we can simply conclude $S_{u}^{-}=\emptyset$ because this implies that $|\partial(u)|=2$ and $\{u,v_{1}\}$ is an edge. Otherwise, if $\partial(v_{1})<\partial(v_{1}+u)$, we can search for $S_{u}^{-}$ as a tipping point for the chain defined by the sequence $v_{1},v_{2},\dots,v_{n-2}$ using Lemma˜C.2 because $\partial(v_{1})<\partial(v_{1}+u)$ and $\partial(v_{1},\dots,v_{n-2})\geq\partial(v_{1},\dots,v_{n-2}+u)$. By the arguments in Lemma˜C.2, there exists $i\in[n-2]$ such that $v_{i}\in N(u)$ and
@@
This implies that $S_{u}^{-}:=L_{i-1}$ satisfies $|\partial(u,S_{u}^{-})|=\lceil\tfrac{|\partial(u)|}{2}\rceil-1$. We can use the same argument to get $S_{u}^{+}$. Moreover, since $S_{u}^{-}$ and $S_{u}^{+}$ are prefix and suffix of the same ordering, and contain no more than half the neighbors of $u$, they must be disjoint. The proof makes two calls to Lemma˜C.2, one for each ordering, with $O(n)$ preprocessing. Each call takes $\widetilde{O}(n)$ time, so the total running time is $\widetilde{O}(n)$. ∎
C.1 Proof of Structural Primitives Lemma
We can now prove the main structural result Theorem˜4.1, which we restate for convenience:
See 4.1 Before we prove the theorem, we first prove a primitive that shows how to identify if a vetrex is isolated using constant queries when $n\geq 4$:
Lemma C.4 (Home Alone).
@@ C.1 Proof of Structural Primitives Lemma
Proof of Theorem˜4.1.
We first check if $\partial(u)=\emptyset$ using Lemma˜C.4. If $\partial(u)=\emptyset$, there are no neighbors of $u$ in $A$ in part (i) and we can certify that $N\cap T=\emptyset$ in part (ii). Assume $\partial(u)\neq\emptyset$ from here on.
In proving part (i), and determining which of $v_{1},\dots,v_{k}\in U\setminus\{u\}$ are neighbors of $u$, we begin by computing the disjoint sets $S_{u}^{-},S_{u}^{+}\subseteq U\setminus\{u\}$ with exactly $\lceil\tfrac{|\partial(u)|}{2}\rceil-1$ neighbors of $u$; this takes $O(\log n)$ queries and $\widetilde{O}(n)$ time due to Lemma˜C.3.
For each $v_{i}$ we choose $S$ to be one of $S_{u}^{-}$ or $S_{u}^{+}$ in which $v_{i}$ does not appear—say $S:=S_{u}^{+}$. Since $|\partial(u,S)|<\tfrac{1}{2}|\partial(u)|$, we know from ˜C.1 that $\text{sign}(|\partial(S)|-|\partial(S+u)|)=-1.$ We then add $v_{i}$ to $S$ and ask for $\text{sign}(|\partial(S+v_{i})|-|\partial(S+v_{i}+u)|)$. Since $\partial(S,u)$ is just one edge short of reaching half the degree of $u$, this sign increases if and only if $v_{i}\in N(u)$. Thus, each test requires one query, and the total query complexity is $k+O(\log n)$. After computing the median sets $S_{u}^{-}$ and $S_{u}^{+}$ in $\widetilde{O}(n)$ time, we partition the vertices $v_{1},\dots,v_{k}$ into two groups based on whether each $v_{i}$ lies in $S_{u}^{-}$ or $S_{u}^{+}$; this takes $O(k)$ time using bit vector lookups. For each group, we fix the corresponding set $S:=S_{u}^{+}$ or $S_{u}^{-}$ and write its bit vector once in $O(n)$ time. Each query $(S+v_{i},S+v_{i}+u)$ is then constructed in $O(1)$ time via bit flipping. The total running time is $\widetilde{O}(n+k)$.
This proves Theorem˜4.1(i).
For part (ii) of the lemma, consider the ordered sequence $T=(v_{1},\dots,v_{t})\subseteq U\setminus\{u\}$ and extend it to a total ordering of $U\setminus\{u\}$ by appending the remaining vertices arbitrarily, forming the sequence $(v_{1},\dots,v_{t},v_{t+1},\dots,v_{n-1})$. Define suffixes $R_{k}:=\{v_{k},\dots,v_{n-1}\}$ for each $k$; by Lemma˜C.3, there exists some $k$ such that
@@
and such a suffix $S:=R_{k}$ can be found using $O(\log n)$ queries and $\widetilde{O}(n)$ time via Lemma˜C.3. To verify that $|\partial(u,T)|>0$, consider the sequence $(v_{t+1},\dots,v_{n-1},v_{1},\dots,v_{t})$ with $T$ placed all the way at the end. The suffix median of this ordering computed using Lemma˜C.3 is the same as $S$ iff $|\partial(u,T)|>0$. Otherwise, define the nested sequence
@@
By Lemma˜C.2, there exists an index $i$ such that the sign of the query $(S_{i-1},S_{i-1}+u)$ is strictly smaller than that of $(S_{i},S_{i}+u)$. The corresponding vertex $v_{i}$ is the first vertex in $T$ adjacent to $u$, and binary search over $i\in[\min(t,k)]$ finds it using $O(\log n)$ queries and $\widetilde{O}(n)$ time. ∎
See 4.3
@@
We assume that $|\partial(u,T)|\leq\left\lfloor\tfrac{1}{2}|\partial(u)|\right\rfloor+1$; otherwise, we split $T$ into $T_{+}:=T\cap S_{u}^{+}$ and $T_{-}:=T\cap S_{u}^{-}$ using the median sets from Lemma˜C.3, and recursively extract edges from each. Both subsets satisfy the same degree bound, and the total query and runtime complexity remains the same.
@@
C.2 Cut sparsifiers using $\widetilde{O}(n)$ queries
Using Lemma˜4.4, we show that we can implement the sparsifier construction from [RSW18, Algorithm 3.4] and the min-cut computation from [RSW18, Algorithm 4.1] using $\widetilde{O}(n/\varepsilon^{2})$ cut comparison queries. The mentioned algorithms and relevant theorems from [RSW18] are mentioned verbatim below:
@@ C.2 Cut sparsifiers using $\widetilde{O}(n)$ queries
Theorem C.5 (Theorem 3.5 from [RSW18]).
@@
Theorem C.6 (Theorem 4.2 from [RSW18]).
@@
Proof of Theorem˜1.5.
For each $0\leq j\leq\log n$, step 2(a) of Algorithm 2 can be implemented using Lemma˜4.4 with $O(|E_{j}|\cdot q_{j}\cdot\log n)+\widetilde{O}(n)$ comparison queries in expectation; which is shown to be $\widetilde{O}(n)$ in the proof of Theorem˜C.5. The runtime bound of the same step is $\widetilde{O}(n^{2})$. Step 2(b)(ii) of Algorithm 2 also takes $\widetilde{O}(n/\varepsilon^{2})$ queries and $\widetilde{O}(n^{2})$ time because the number of edges sampled and added to $H$ is shown to be $\widetilde{O}(n/\varepsilon^{2})$ (see proof of Theorem 4.2, [RSW18]) and this sub-sampling can be done again using Lemma˜4.4. In Algorithm 3, since step (1) is only used to compare the singleton cuts to output the minimum cut seen by the algorithm, we can implement this using comparison in O(n){O}(n) queries and time. In step (3), all the approximate cuts can be computed in $\widetilde{O}(n^{2})$ time using [HHS24]. In step (4), all the edges of $G^{\prime}$ can be learned using Lemma˜4.4 with $p=1$. The proof of theorem˜C.6 shows that the number of edges in $G^{\prime}$ is at most $O(n)$ with high probability. This implies that all the steps in the algorithms take $\widetilde{O}(n/\varepsilon^{2})$ queries and $\widetilde{O}(n^{2})$ time. There are at most $\log n$ steps in any algorithm which proves the desired bounds. ∎
C.3 Weighted Min-Cut: Special Cases and Stronger Oracles
In the weighted setting, where the edges of the graph $G=(V,E)$ have unknown non-negative weights, the techniques developed for the unweighted case (where weights are 0 or $1$; see Section˜4) no longer apply. Even when the weights are restricted to integers in $[0,B]$, little is known.
C.3.1 Weighted regular graphs

### v2

@@ A.1 The Certification Problem
Before addressing the optimization problem, we ask whether it is even possible to certify optimality efficiently. Suppose we wish to prove that some point $y\in\mathcal{P}$ is the optimal solution, i.e., $y\in\arg\min_{x\in\mathcal{P}}\langle w^{*},x\rangle$, or equivalently, that
@@ A.3 Beyond the Hypercube
Appendix B Omitted proofs from Section˜3
Proof of Lemma˜3.4.
For the first part, observe that $C_{\pi}$ is the intersection of N1N-1 closed half-spaces, and is thus a convex polyhedron.
Assume $C_{\pi}$ is full-dimensional. Then its interior, $\operatorname{int}(C_{\pi})$, is non-empty. For any $w\in\operatorname{int}(C_{\pi})$, strict inequalities $\langle w,x_{\pi(1)}\rangle<\langle w,x_{\pi(2)}\rangle<\dots<\langle w,x_{\pi(N)}\rangle$ hold. Consequently, $\operatorname{int}(C_{\pi})$ does not intersect any hyperplane in $\mathcal{A}$. Since $C_{\pi}$ is convex, it is connected. To show $C_{\pi}$ is a cell, we check maximality. Suppose $S\subseteq\mathbb{R}^{d}$ is a connected region such that $C_{\pi}\subseteq S$ and $\operatorname{int}(S)\cap\mathcal{A}=\emptyset$. Since $S$ is connected and its interior avoids all hyperplanes $h_{\{i,j\}}$, the sign of $\langle w,x_{i}-x_{j}\rangle$ must be constant for all $w\in\operatorname{int}(S)$. Because $\operatorname{int}(C_{\pi})\subset\operatorname{int}(S)$, this sign pattern is determined by $\pi$. Thus, every $w\in S$ must satisfy the constraints defining $C_{\pi}$, implying $S\subseteq C_{\pi}$. Therefore, $C_{\pi}$ is a maximal connected region whose interior avoids $\mathcal{A}$, making it a cell.
Conversely, let $C$ be any cell of the arrangement $\mathcal{A}$. Since cells are full-dimensional, $\operatorname{int}(C)$ is non-empty. Pick any $w\in\operatorname{int}(C)$; the values $\{\langle w,x_{k}\rangle\}_{k=1}^{N}$ are distinct and induce a permutation $\pi\in S_{N}$. Since $C$ is connected and $\operatorname{int}(C)\cap\mathcal{A}=\emptyset$, the ordering $\pi$ is constant over $\operatorname{int}(C)$. Thus $C\subseteq C_{\pi}$. Since $C$ is a maximal region with this property and $C_{\pi}$ is a region satisfying the property (as shown above), it must be that $C=C_{\pi}$.
For the second part regarding partial orders, $C_{\sigma}$ is the intersection of the half-spaces $H_{j,i}$ for all pairs $i\preceq_{\sigma}j$, which implies it is a convex polyhedron. To prove the decomposition, consider any $w\in C_{\sigma}$. The scalar values $\{\langle w,x_{k}\rangle\}_{k=1}^{N}$ induce a total ordering. Since $w$ satisfies the constraints of $\sigma$, this total ordering is a linear extension of $\sigma$. Thus, there exists some $\pi\in\mathcal{L}(\sigma)$ such that $w\in C_{\pi}$. Conversely, if $w\in C_{\pi}$ for some $\pi\in\mathcal{L}(\sigma)$, then by transitivity, $w$ satisfies $\langle w,x_{i}\rangle\leq\langle w,x_{j}\rangle$ for all $i\preceq_{\sigma}j$, implying $w\in C_{\sigma}$. Therefore, $C_{\sigma}=\bigcup_{\pi\in\mathcal{L}(\sigma)}C_{\pi}$. ∎
Proof of Lemma˜3.5.
By definition, the partial order $\sigma^{*}$ contains exactly those relations $i\preceq_{\sigma^{*}}j$ for which $\langle w^{*},x_{j}-x_{i}\rangle>0$. Since $C_{\sigma^{*}}$ is defined by the intersection of half-spaces $\langle w,x_{j}-x_{i}\rangle\geq 0$ for all $i\preceq_{\sigma^{*}}j$, and $w^{*}$ satisfies all these inequalities strictly, $w^{*}$ must lie in the interior of $C_{\sigma^{*}}$.
For the second part, we prove the contrapositive: $\langle w^{*},x_{i}-x_{j}\rangle>0\implies\langle\widehat{w},x_{i}-x_{j}\rangle>0$. Suppose $\langle w^{*},x_{i}-x_{j}\rangle>0$. By the definition of $\sigma^{*}$, this strict inequality implies the relation $j\preceq_{\sigma^{*}}i$. Consequently, every point $w\in C_{\sigma^{*}}$ must satisfy $\langle w,x_{i}-x_{j}\rangle\geq 0$. Since $\widehat{w}$ is in the interior of $C_{\sigma^{*}}$, it must strictly satisfy the inequalities implied by the partial order. Therefore, $\langle\widehat{w},x_{i}-x_{j}\rangle>0$. ∎
Proof of ˜3.6.
By the definition of the CCO output for $\widehat{w}$, the sequence is non-decreasing with respect to $\widehat{w}$, meaning $\langle\widehat{w},y_{i}-y_{i+1}\rangle\leq 0$ for all $1\leq i\leq k-1$. Applying Lemma˜3.5, the implication
$\langle\widehat{w},x_{u}-x_{v}\rangle\leq 0\implies\langle w^{*},x_{u}-x_{v}\rangle\leq 0$
holds for all pairs from $\mathcal{P}$. Therefore, we have $\langle w^{*},y_{i}-y_{i+1}\rangle\leq 0$ for all $i$. By transitivity, $\langle w^{*},y_{i}-y_{j}\rangle\leq 0$ for all $1\leq i\leq j\leq k$.
Finally, using the property that $\mathcal{P}\subseteq y_{1}+\operatorname{env}(y_{1},\dots,y_{k})$, and observing that $w^{*}$ makes a non-negative inner product with the generators of this envelope (as implied by the sorted order), it follows that $y_{1}\in\operatorname*{argmin}_{x\in\mathcal{P}}\langle w^{*},x\rangle$. ∎
Appendix C Omitted proofs from Section˜4
Proof of Lemma˜4.6.
@@ Appendix C Omitted proofs from Section˜4
Appendix D Omitted proofs from Section˜5
@@
Lemma D.2 (tipping point).
@@
By ˜D.1, the $\textnormal{sign}(|\partial(S)|-|\partial(S+u)|)$ is monotonic in $S$, since the quantity $|\partial(u,S)|$ increases as $S$ grows. Therefore, if the sign changes between $S_{0}$ and $S_{t}$, then by monotonicity, it must strictly increase at some index $i$, which proves (22). Moreover, using the equality in (21) again, (22) can be rewritten as
@@
Lemma D.3 (median sets).
@@
Let $L_{j}=\{v_{1},v_{2},\dots\dots,v_{j}\}$ and $R_{j}=\{v_{j},v_{j+1},\dots,v_{n-1}\}$ for $j\in[n-1]$. First, check if $\partial(v_{1})>\partial(v_{1}+u)$ in which case, we can conclude $S_{u}^{-}=\emptyset,S_{u}^{+}=\{v_{2},v_{3},\dots,v_{n-1}\}$ because $\{u,v_{1}\}$ has to be the only edge using ˜D.1. Similarly for $v_{n-1}$. So from now on, assume that $\partial(v_{1})\leq\partial(v_{1}+u)$ and $\partial(v_{n-1})\leq\partial(v_{n-1}+u)$. Equivalently, this can also be written as $\partial(R_{2})\geq\partial(R_{2}+u),\partial(L_{n-2})\geq\partial(L_{n-2}+u)$.
In what follows, we will first describe how to compute $S_{u}^{-}$ and then use a similar argument to compute $S_{u}^{+}$. If $\partial(v_{1})=\partial(v_{1}+u)$, then again we can simply conclude $S_{u}^{-}=\emptyset$ because this implies that $|\partial(u)|=2$ and $\{u,v_{1}\}$ is an edge. Otherwise, if $\partial(v_{1})<\partial(v_{1}+u)$, we can search for $S_{u}^{-}$ as a tipping point for the chain defined by the sequence $v_{1},v_{2},\dots,v_{n-2}$ using Lemma˜D.2 because $\partial(v_{1})<\partial(v_{1}+u)$ and $\partial(v_{1},\dots,v_{n-2})\geq\partial(v_{1},\dots,v_{n-2}+u)$. By the arguments in Lemma˜D.2, there exists $i\in[n-2]$ such that $v_{i}\in N(u)$ and
@@
This implies that $S_{u}^{-}:=L_{i-1}$ satisfies $|\partial(u,S_{u}^{-})|=\lceil\tfrac{|\partial(u)|}{2}\rceil-1$. We can use the same argument to get $S_{u}^{+}$. Moreover, since $S_{u}^{-}$ and $S_{u}^{+}$ are prefix and suffix of the same ordering, and contain no more than half the neighbors of $u$, they must be disjoint. The proof makes two calls to Lemma˜D.2, one for each ordering, with $O(n)$ preprocessing. Each call takes $\widetilde{O}(n)$ time, so the total running time is $\widetilde{O}(n)$. ∎
D.1 Proof of Structural Primitives Lemma
We can now prove the main structural result Theorem˜5.1, which we restate for convenience:
See 5.1 Before we prove the theorem, we first prove a primitive that shows how to identify if a vetrex is isolated using constant queries when $n\geq 4$:
Lemma D.4 (Home Alone).
@@ D.1 Proof of Structural Primitives Lemma
Proof of Theorem˜5.1.
We first check if $\partial(u)=\emptyset$ using Lemma˜D.4. If $\partial(u)=\emptyset$, there are no neighbors of $u$ in $A$ in part (i) and we can certify that $N\cap T=\emptyset$ in part (ii). Assume $\partial(u)\neq\emptyset$ from here on.
In proving part (i), and determining which of $v_{1},\dots,v_{k}\in U\setminus\{u\}$ are neighbors of $u$, we begin by computing the disjoint sets $S_{u}^{-},S_{u}^{+}\subseteq U\setminus\{u\}$ with exactly $\lceil\tfrac{|\partial(u)|}{2}\rceil-1$ neighbors of $u$; this takes $O(\log n)$ queries and $\widetilde{O}(n)$ time due to Lemma˜D.3.
For each $v_{i}$ we choose $S$ to be one of $S_{u}^{-}$ or $S_{u}^{+}$ in which $v_{i}$ does not appear—say $S:=S_{u}^{+}$. Since $|\partial(u,S)|<\tfrac{1}{2}|\partial(u)|$, we know from ˜D.1 that $\text{sign}(|\partial(S)|-|\partial(S+u)|)=-1.$ We then add $v_{i}$ to $S$ and ask for $\text{sign}(|\partial(S+v_{i})|-|\partial(S+v_{i}+u)|)$. Since $\partial(S,u)$ is just one edge short of reaching half the degree of $u$, this sign increases if and only if $v_{i}\in N(u)$. Thus, each test requires one query, and the total query complexity is $k+O(\log n)$. After computing the median sets $S_{u}^{-}$ and $S_{u}^{+}$ in $\widetilde{O}(n)$ time, we partition the vertices $v_{1},\dots,v_{k}$ into two groups based on whether each $v_{i}$ lies in $S_{u}^{-}$ or $S_{u}^{+}$; this takes $O(k)$ time using bit vector lookups. For each group, we fix the corresponding set $S:=S_{u}^{+}$ or $S_{u}^{-}$ and write its bit vector once in $O(n)$ time. Each query $(S+v_{i},S+v_{i}+u)$ is then constructed in $O(1)$ time via bit flipping. The total running time is $\widetilde{O}(n+k)$.
This proves Theorem˜5.1(i).
For part (ii) of the lemma, consider the ordered sequence $T=(v_{1},\dots,v_{t})\subseteq U\setminus\{u\}$ and extend it to a total ordering of $U\setminus\{u\}$ by appending the remaining vertices arbitrarily, forming the sequence $(v_{1},\dots,v_{t},v_{t+1},\dots,v_{n-1})$. Define suffixes $R_{k}:=\{v_{k},\dots,v_{n-1}\}$ for each $k$; by Lemma˜D.3, there exists some $k$ such that
@@
and such a suffix $S:=R_{k}$ can be found using $O(\log n)$ queries and $\widetilde{O}(n)$ time via Lemma˜D.3. To verify that $|\partial(u,T)|>0$, consider the sequence $(v_{t+1},\dots,v_{n-1},v_{1},\dots,v_{t})$ with $T$ placed all the way at the end. The suffix median of this ordering computed using Lemma˜D.3 is the same as $S$ iff $|\partial(u,T)|>0$. Otherwise, define the nested sequence
@@
By Lemma˜D.2, there exists an index $i$ such that the sign of the query $(S_{i-1},S_{i-1}+u)$ is strictly smaller than that of $(S_{i},S_{i}+u)$. The corresponding vertex $v_{i}$ is the first vertex in $T$ adjacent to $u$, and binary search over $i\in[\min(t,k)]$ finds it using $O(\log n)$ queries and $\widetilde{O}(n)$ time. ∎
See 5.3
@@
We assume that $|\partial(u,T)|\leq\left\lfloor\tfrac{1}{2}|\partial(u)|\right\rfloor+1$; otherwise, we split $T$ into $T_{+}:=T\cap S_{u}^{+}$ and $T_{-}:=T\cap S_{u}^{-}$ using the median sets from Lemma˜D.3, and recursively extract edges from each. Both subsets satisfy the same degree bound, and the total query and runtime complexity remains the same.
@@
D.2 Cut sparsifiers using $\widetilde{O}(n)$ queries
Using Lemma˜5.4, we show that we can implement the sparsifier construction from [RSW18, Algorithm 3.4] and the min-cut computation from [RSW18, Algorithm 4.1] using $\widetilde{O}(n/\varepsilon^{2})$ cut comparison queries. The mentioned algorithms and relevant theorems from [RSW18] are mentioned verbatim below:
@@ D.2 Cut sparsifiers using $\widetilde{O}(n)$ queries
Theorem D.5 (Theorem 3.5 from [RSW18]).
@@
Theorem D.6 (Theorem 4.2 from [RSW18]).
@@
Proof of Theorem˜1.6.
For each $0\leq j\leq\log n$, step 2(a) of Algorithm 5 can be implemented using Lemma˜5.4 with $O(|E_{j}|\cdot q_{j}\cdot\log n)+\widetilde{O}(n)$ comparison queries in expectation; which is shown to be $\widetilde{O}(n)$ in the proof of Theorem˜D.5. The runtime bound of the same step is $\widetilde{O}(n^{2})$. Step 2(b)(ii) of Algorithm 5 also takes $\widetilde{O}(n/\varepsilon^{2})$ queries and $\widetilde{O}(n^{2})$ time because the number of edges sampled and added to $H$ is shown to be $\widetilde{O}(n/\varepsilon^{2})$ (see proof of Theorem 4.2, [RSW18]) and this sub-sampling can be done again using Lemma˜5.4. In Algorithm 6, since step (1) is only used to compare the singleton cuts to output the minimum cut seen by the algorithm, we can implement this using comparison in O(n){O}(n) queries and time. In step (3), all the approximate cuts can be computed in $\widetilde{O}(n^{2})$ time using [HHS24]. In step (4), all the edges of $G^{\prime}$ can be learned using Lemma˜5.4 with $p=1$. The proof of theorem˜D.6 shows that the number of edges in $G^{\prime}$ is at most $O(n)$ with high probability. This implies that all the steps in the algorithms take $\widetilde{O}(n/\varepsilon^{2})$ queries and $\widetilde{O}(n^{2})$ time. There are at most $\log n$ steps in any algorithm which proves the desired bounds. ∎
D.3 Weighted Min-Cut: Special Cases and Stronger Oracles
In the weighted setting, where the edges of the graph $G=(V,E)$ have unknown non-negative weights, the techniques developed for the unweighted case (where weights are 0 or $1$; see Section˜5) no longer apply. Even when the weights are restricted to integers in $[0,B]$, little is known.
D.3.1 Weighted regular graphs